OpenStudy (anonymous):

Juan drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 12 hours. When Juan drove home, there was no traffic and the trip only took 8 hours. If his average rate was 20 miles per hour faster on the trip home, how far away does Juan live from the mountains? Do not do any rounding.

OpenStudy (anonymous):

There is something wrong with the problem statement. It should not take a longer period of time to return when the speed is higher and the distances are equal.

OpenStudy (anonymous):

I corrected it, sorry my bad lol

OpenStudy (anonymous):

r * t = d Let x be the speed on the way to the mountains. Solve the following for x: 12 x = 8 (x + 20) x = 40 mph 12 x = 12 * 40 = 480 miles